A 725 N man stands in the middle of a frozen pond of radius 9.0 m. He is unable to get to the other side because of a lack of friction between his shoes and the ice. To overcome this difficulty, he throws his 1.2 kg physics textbook horizontally toward the north shore at a speed of 9.0 m/s. How long does it take him to reach the south shore?
s
Answer #1

apply momentum conservation

m*v = m1*v1

velocity of man = 1.2 * 9 * 9.81/ (725 ) =

time = 9 / 0.1461 = 61.60 sec